OpenStudy (anonymous):

Determine the midpoint given the following points . A(4,2) B (-8,18) ??

OpenStudy (anonymous):

use the midpoint formula: (average of the x's, average of the y's)

OpenStudy (anonymous):

Here's a visual representation of what dpalnc said: \[(\frac{x_{1} + x_{2}}{2}, \frac{y_{1} + y_{2}}{2})\]

OpenStudy (anonymous):

\[\text{mid}(A,B)=\left(\frac{\Delta x}{2}, \frac{\Delta y}{2}\right)\]

OpenStudy (anonymous):

errr.... let \(\Delta x=x_2+x_1\) and \(\Delta y=y_2+y_1\)..

OpenStudy (anonymous):

I came up with (-2,10) Please say im correct ?

OpenStudy (anonymous):

You are. Good job.

OpenStudy (anonymous):

Thakss :)

OpenStudy (anonymous):

You're welcome. Have a good day now.

OpenStudy (anonymous):

You too . & thanks for your help today. I learned alot

OpenStudy (anonymous):

Yay! Mission accomplished.

OpenStudy (anonymous):

lol... "please say i'm correct"... :)